# arXiv:2603.01359v1[eess.SY] 02 Mar 2026

### v1

@@ Abstract
Over the last decades, progress in modal analysis has enabled increasingly routine use of modal parameters, including those extracted from in-situ measurements, for applications such as structural health monitoring and finite element model updating. For output-only identification, or Operational Modal Analysis (OMA), widely adopted approaches include Stochastic Subspace Identification (SSI) methods and the Natural Excitation Technique combined with the Eigensystem Realization Algorithm (NExT-ERA). Nevertheless, SSI-based techniques may become cumbersome on large systems, while NExT-ERA fitting can struggle when measurements are contaminated by noise. To alleviate these, this work investigates an OMA frequency-domain formulation for aeronautical structures by coupling the Loewner Framework (LF) with NExT, yielding the proposed NExT-LF method. The method exploits the computational efficiency of LF together with the impulse response function retrieval enabled by NExT. NExT-LF is assessed on two experimental benchmarks: the eXperimental BeaRDS 2 high-aspect-ratio wing main spar and an Airbus Helicopters H135 bearingless main rotor blade. The identified modal parameters are compared against available experimental references and results obtained via SSI with Canonical Variate Analysis and NExT-ERA. The results show that the modes identified by NExT-LF correlate well with benchmark data, particularly for high-amplitude tests and in the low-frequency range.
@@ 1 Introduction
Modal analysis is a key tool for characterising the structural dynamics of engineering systems Ewins (2000). It can be performed in an input-controlled experimental setting or by measuring a structure in situ, either during operation or at rest (Ambient Vibration Testing – AVT), without the application of an external controlled input. The former is referred to as Experimental Modal Analysis (EMA), whereas the latter is known as Operational Modal Analysis (OMA) Brincker and Ventura (2015). In OMA, which is the focus of this work, the objective is to extract modal parameters Reynders (2012)—natural frequencies ($\omega_{n}$), damping ratios ($\zeta_{n}$), and mode shapes ($\bm{\phi}_{n}$)—to describe the dynamic behaviour of a structure for purposes including Finite Element Model (FEM) updating Zanotti Fragonara et al. (2017) and Structural Health Monitoring (SHM) Cadoret et al. (2025). In aeronautical and aerospace engineering, OMA is employed in a broad range of applications. For instance, Jelicic et al. (2014) uses OMA results to track flutter stability in real time for a wind-tunnel wing model, while Sinske et al. (2017) reports in-flight aeroelastic identification for the German Aerospace Center (DLR) HALO research aircraft. OMA has also been applied during launch-vehicle rollout operations, as reported for Artemis I in Akers et al. (2025), and to extract modal characteristics from in-flight data for launch vehicles and missiles in Eugeni et al. (2018) and James et al. (1994), respectively. Rotorcraft applications are also well established: Ameri et al. (2013) applies OMA to Ground Vibration Testing (GVT) data from a Westland Lynx Mark 7 helicopter airframe, while Sibille et al. (2023) and Agneni et al. (2010) identify time-expired helicopter blades from an Airbus Helicopters H135 and an Agusta-Bell AB 204, respectively.
@@
Traditional OMA techniques include the Natural Excitation Technique (NExT) paired with the Eigensystem Realization Algorithm (ERA), commonly denoted as NExT-ERA James III et al. (1993), Stochastic Subspace Identification with Canonical Variate Analysis (SSI-CVA) Van Overschee and De Moor (1996), and Frequency Domain Decomposition (FDD) Brincker et al. (2001). NExT-ERA and SSI-CVA are time-domain approaches, whereas FDD operates directly in the frequency domain. Recent developments have aimed at improving robustness to measurement noise and supporting consistent mode selection. For example, O’Connell and Rogers (2024); O’Connell et al. (2025) propose a robust COVariance-driven SSI (SSI-COV) formulation incorporating a probabilistic criterion to reduce the likelihood that outliers are misinterpreted as physical modes. With a related objective, Amador and Brincker (2024) introduces a subspace-based poly-reference complex frequency approach to reduce uncertainties associated with the null space of the system matrices used in the identification process, by adopting a subspace-driven parametric formulation and estimating modal properties via eigensystem realisations across selected subspaces. Another SSI-based development couples SSI-COV with Fuzzy C-Means (SSI-FCM) clustering to reject spurious modes in stabilisation diagrams, and has been demonstrated on a train bogie and a five-storey experimental structure in Lv et al. (2026). Furthermore, the Multi-Reference Transmissibility Complex Frequency method is proposed in Tarinejad and Amanzad (2026) to address bias errors related to the application of the Fast Fourier Transform (FFT) to non-stationary signals, although validation is limited to numerical systems. Finally, NExT combined with Dynamic Mode Decomposition (DMD) is introduced for SHM of aeronautical structures in Wu et al. (2026), building on the DMD formulation for experimental modal analysis in Saito and Kuno (2020), which in turn extends the original DMD framework for extracting coherent features in fluid flows Schimid (2010).
@@
In parallel, automated OMA frameworks (AOMA) have been proposed to streamline the identification process, including machine learning-based strategies Mugnaini et al. (2022), hierarchical clustering Neu et al. (2017), k-means clustering de Almeida Cardoso et al. (2018), and Density-Based Spatial Clustering of Applications with Noise (DBSCAN) Sibille et al. (2023). AOMA is outside the scope of this work, and the interested reader is referred to Mostafaei and Ghamami (2025) for a recent review.
@@
Within this broader context, the frequency-domain Loewner Framework (LF) has been considered for EMA Dessena et al. (2023) and subsequently extended to output-only identification in civil engineering by pairing it with NExT, yielding the NExT-LF approach Dessena et al. (2025). The aim of this coupling is to exploit the computational efficiency of the LF fitting processs while retaining the output-only capability provided by NExT. These characteristics make NExT-LF attractive for aeronautical applications, where datasets can be large and measurements may be affected by operational noise. To address this gap, the present work investigates NExT-LF on two aeronautically relevant experimental datasets: the eXperimental BeaRDS 2 (XB-2) flexible wing spar model Dessena et al. (2022) and an Airbus Helicopters H135 Bearingless Main Rotor (BMR) blade Weber et al. (2021). The two datasets are selected due to their different levels of complexity in terms of system dimensions (1.45 m vs. 4.99 m span) and test conditions (shaker excitation vs. AVT). The identification accuracy of NExT-LF is evaluated against concurrent identifications obtained via SSI-CVA and NExT-ERA, as well as against available benchmark results. The overall objective is to assess whether the accuracy achieved by NExT-LF is suitable for aeronautical applications. To achieve this aim, the remainder of the paper is structured as follows:
@@
  • Section˜2 (Methodology) introduces the LF identification procedure, the output-only enabling method (NExT), and their coupling into NExT-LF;
@@ 2.1 Loewner Framework
The LF algorithm has previously been adopted for the modelling of multi-port electrical systems Lefteriu and Antoulas (2009) and, in the field of aeroservoelastic modelling, for aerodynamic model order reduction Quero et al. (2019). Subsequently, the first and second authors employed the LF to identify modal parameters in SIMO mechanical systems Dessena et al. (2023), demonstrated its computational efficiency Dessena et al. (2023), and examined its robustness to measurement noise for SHM applications Dessena et al. (2024). More recent work has extended the LF to enable the extraction of modal parameters from multi-input multi-output Dessena and Civera (2025); Dessena et al. (2026) and output-only systems Dessena et al. (2025). The present study considers this latter output-only formulation.
To introduce the LF properly, we first define the Loewner matrix $\bm{\mathbb{L}}$}: Let ($\mu_{j}$${v}_{j}$), j=1j=1,…,$q$ denote a row-wise collection of complex pairs and ($\lambda_{i}$${w}_{j}$), i=1i=1,…,$k$ a column-wise collection of complex pairs, with all points $\lambda_{i}$ and $\mu_{j}$ mutually distinct. The matrix $\bm{\bm{\mathbb{L}}}$, commonly termed the divided-differences $\bm{\mathbb{L}}$} matrix, is then given by:
@@
Karl Löwner first linked $\bm{\mathbb{L}}$ to rational interpolation, a setting that is frequently described as Cauchy interpolation Löwner (1934). Through this link, interpolants may be characterised via the determinants of suitable submatrices of $\bm{\mathbb{L}}$. Moreover, as shown in Antoulas et al. (2017); Mayo and Antoulas (2007), one can obtain rational interpolants directly from $\bm{\mathbb{L}}$. In the present work, the formulation adopted is the Loewner pencil, defined by the pair of matrices $\bm{\mathbb{L}}$ and $\bm{\mathbb{L}}_{s}$. The latter, $\bm{\mathbb{L}}_{s}$, is the shifted Loewner matrix, introduced later on.
@@
The Laplace-domain transfer function $\bm{H}(s)$ corresponding to $\bm{\Sigma}$ may be expressed as a $p\times m$ rational matrix function, provided that the pencil $\bm{A}-\lambda\bm{E}$ is regular, i.e., $\bm{A}-\lambda\bm{E}$ is non-singular for at least one finite $\lambda\in\mathbb{C}$:
@@
Let us consider the general setting of tangential interpolation, which is commonly described as rational interpolation along tangential directions Kramer and Gugercin (2016). The associated right interpolation dataset is defined as:
@@
The values $\lambda_{i}$ and $\mu_{j}$ denote the sampling points at which $\bm{H}(s)$ is queried, which in the present context coincide with the chosen frequency bins. The vectors $\bm{r}_{i}$ and $\bm{l}_{j}$ specify the right and left tangential directions and are often drawn at random in practical implementations Quero et al. (2019). The quantities $\bm{w}_{i}$ and $\bm{v}_{j}$ are the corresponding tangential measurements. The rational interpolation task is then satisfied by enforcing consistency between $\bm{w}_{i}$, $\bm{v}_{j}$ and the transfer function $\bm{H}$ associated with the realisation $\bm{\Sigma}$ in Equation˜2:
@@
A realisation of minimal order is attainable only when the underlying system is both controllable and observable. Under the assumption that the available samples originate from a system whose transfer function is given by Equation˜13, the generalised tangential observability matrix $\mathcal{O}_{v}$ and the generalised tangential controllability matrix $\mathcal{R}_{\rho}$ are introduced in Lefteriu and Antoulas (2010). It then follows that Equation˜9 and Equation˜11 may be recast in the form:
@@
The expression in Equation˜20, which specifies the Loewner realisation in the presence of redundant data, is adopted as the reference formulation in the remainder of this work. The complete step-by-step account procedure is provided in Mayo and Antoulas (2007); Antoulas et al. (2017). Finally, the modal parameters of the system may be obtained by performing an eigenanalysis of the matrices $\bm{A}$ and $\bm{C}$ appearing in Equation˜20. In practical modal analysis, this procedure is repeated over a range of model orders $k$, selected to capture the minimum set of modes of interest while identifying the highest order for which these modes remain stably estimated. From a practical standpoint, the frequency domain data, usually an FRF, $\bm{H}_{real}(s)$ is fed to the LF, which then breaks down the data into two subsets, the left and right data, so that the system matrices of the transfer function in Equation˜13 can be retrieved.
@@ 2.2 Natural Excitation Technique
NExT enables modal identification from output-only measurements by exploiting the statistical structure of ambient (unmeasured) excitations. In many operational settings, the applied forces cannot be recorded reliably (e.g. wind, traffic, or in-service aerodynamic loads), which prevents the direct construction of FRFs from measured inputs and outputs. NExT circumvents this limitation by assuming the unknown excitation as a random process with sufficiently broad frequency content to excite the modes of interest, and by using output correlations as impulse-response-type quantities James III et al. (1993), the system IRFs can be obtained.
@@
where $\tau$ is the time lag. For finite records, (22) is estimated from sampled data. Using $N$ samples $y_{i}[k]$ and $y_{j}[k]$, one common unbiased estimator for non-negative lags is
@@
When the ambient excitation is broadband and the system is LTI, the correlation functions can be written as a superposition of decaying sinusoids whose damped frequencies and decay rates coincide with those of the structural modes James III et al. (1993). In practice, this means that $\widehat{R}_{y_{i}y_{j}}[\ell]$ (with $\ell$ denoting the discrete lag index, i.e. $\tau=\ell\Delta t$ for sampling period $\Delta t$) has the same functional form as a free-decay (IRF) for modal identification purposes, up to an unknown scaling and sensor-pair-dependent phase. A convenient representation is:
@@
where $n_{m}$ is the number of participating modes, $\omega_{r}$ and $\zeta_{r}$ are the natural frequency and damping ratio of mode $r$, and $A_{r,ij}$ and $\theta_{r,ij}$ are constants that depend on the sensor pair and on the (unknown) spatial distribution of the excitation.
Given $p$ output channels, NExT forms a set of impulse-response-like sequences by selecting a reference channel $j$ and computing $\widehat{R}_{y_{i}y_{j}}[\ell]$ for $i=1,\dots,p$. The resulting collection $\{\widehat{R}_{y_{1}y_{j}},\widehat{R}_{y_{2}y_{j}},\dots,\widehat{R}_{y_{p}y_{j}}\}$ is interpreted as a multiple-output IRF associated with the chosen reference. An equivalent frequency-domain route, not considered in this work, computes cross-spectral densities and then obtains correlations via an inverse Fourier transform. Both approaches are consistent provided the stationarity assumptions are satisfied and the same normalisation conventions are used.

### v2

@@ Abstract
Over the last decades, progress in modal analysis has enabled increasingly routine use of modal parameters for applications such as structural health monitoring and finite element model updating. For output-only identification, or Operational Modal Analysis (OMA), widely adopted approaches include Stochastic Subspace Identification (SSI) methods and the Natural Excitation Technique combined with the Eigensystem Realization Algorithm (NExT-ERA). Nevertheless, SSI-based techniques may become cumbersome on large systems, while NExT-ERA fitting can struggle when measurements are contaminated by noise. To alleviate these, this work investigates an OMA frequency-domain formulation for aeronautical structures by coupling the Loewner Framework (LF) with NExT, yielding the proposed NExT-LF method. The method exploits the computational efficiency of LF, due to the effectiveness of tangential interpolation, together with the impulse response function retrieval enabled by NExT. NExT-LF is assessed on two experimental benchmarks: the eXperimental BeaRDS 2 high-aspect-ratio wing main spar and an Airbus Helicopters H135 bearingless main rotor blade. The identified modal parameters are compared against available experimental references and results obtained via SSI with Canonical Variate Analysis and NExT-ERA. The results show that the modes identified by NExT-LF correlate well with benchmark data, particularly for high-amplitude tests and in the low-frequency range.
@@ 1 Introduction
Modal analysis is a key tool for characterising the structural dynamics of engineering systems Ewins (2000). It can be performed in an input-controlled experimental setting or by measuring a structure in situ, either during operation or at rest (Ambient Vibration Testing – AVT), without the application of an external controlled input. The former is referred to as Experimental Modal Analysis (EMA), while the latter is known as Operational Modal Analysis (OMA) Rainieri and Fabbrocino (2014). In OMA, which is the focus of this work, the objective is to extract modal parameters Reynders (2012) – natural frequencies ($\omega_{n}$), damping ratios ($\zeta_{n}$), and mode shapes ($\bm{\phi}_{n}$) – to describe the dynamic behaviour of a structure for purposes including Finite Element Model (FEM) updating Fragonara et al. (2017) and Structural Health Monitoring (SHM) Cadoret et al. (2025). In aeronautical and aerospace engineering, OMA is employed in a wide range of applications. For instance, Jelicic et al. (2014) uses OMA results to track flutter stability in real time for a wind-tunnel wing model, while Sinske et al. (2017) reports in-flight aeroelastic identification for the German Aerospace Center (DLR) HALO research aircraft. OMA has also been applied during launch vehicles rollout operations, as reported for Artemis I in Akers et al. (2025), and to extract modal characteristics from in-flight data for launch vehicles and missiles in Eugeni et al. (2018) and James et al. (1994), respectively. Rotorcraft applications are also well established: Ameri et al. (2013) applies OMA to Ground Vibration Testing (GVT) data from a Westland Lynx Mark 7 helicopter airframe, while Sibille et al. (2023) and Agneni et al. (2010) identify time-expired helicopter blades from an Airbus Helicopters H135 and an Agusta-Bell AB 204, respectively.
@@
Traditional OMA techniques include the Natural Excitation Technique (NExT) paired with the Eigensystem Realization Algorithm (ERA), commonly denoted NExT-ERA III et al. (1993), Stochastic Subspace Identification with Canonical Variate Analysis (SSI-CVA) Overschee and Moor (1996), and Frequency Domain Decomposition (FDD) Brincker et al. (2001). NExT-ERA and SSI-CVA are time-domain approaches, whereas FDD operates directly in the frequency domain. Recent developments have aimed to improve the robustness to measurement noise and support consistent mode selection. For example, O’Connell and Rogers (2024); O’Connell et al. (2025) propose a robust COVariance-driven SSI (SSI-COV) formulation that incorporates a probabilistic criterion to reduce the likelihood that outliers are misinterpreted as physical modes. With a related objective, Amador and Brincker (2024) introduces a subspace-based Poly-reference Complex Frequency approach to reduce uncertainties associated with the null space of the system matrices used in the identification process, by adopting a subspace-driven parametric formulation and estimating modal properties via eigensystem realisations across selected subspaces. Another SSI-based development couples SSI-COV with Fuzzy C-Means (SSI-FCM) clustering to reject spurious modes in stabilisation diagrams, and has been demonstrated on applications to a train bogie and a five-storey experimental structure in Lv et al. (2026). Furthermore, the Multi-Reference Transmissibility Complex Frequency method is proposed in Tarinejad and Amanzad (2026) to address bias errors related to the application of the Fast Fourier Transform (FFT) to non-stationary signals, although validation is limited to numerical systems. Finally, NExT combined with Dynamic Mode Decomposition (DMD) is introduced for OMA on aeronautical structures in Wu et al. (2026), based on the DMD formulation for experimental modal analysis in Saito and Kuno (2020); Wu et al. (2025), which in turn extend the original DMD framework for extracting coherent features in fluid flows Schimid (2010).
@@
In parallel, automated OMA frameworks (AOMA) have been proposed to streamline the identification process, including machine learning-based strategies Mugnaini et al. (2022), hierarchical clustering Neu et al. (2017), k-means clustering de Almeida Cardoso et al. (2018), and Density-Based Spatial Clustering of Applications with Noise (DBSCAN) Sibille et al. (2023). The interested reader is referred to Mostafaei and Ghamami (2025) for a recent review on AOMA applications.
@@
Within this broader context, the frequency-domain Loewner Framework (LF) has been considered for EMA Dessena et al. (2023) and subsequently extended to output-only identification in civil engineering by pairing it with NExT, yielding the NExT-LF approach Dessena et al. (2025). The aim of this coupling is to exploit the computational efficiency of the LF fitting process while retaining the output-only capability provided by NExT. These characteristics make NExT-LF attractive for aeronautical applications, where datasets can be large and measurements may be affected by operational noise. To address this gap, the present work investigates NExT-LF on applications to two aeronautically relevant experimental datasets: the eXperimental BeaRDS 2 (XB-2) flexible wing spar model Dessena et al. (2022) and an Airbus Helicopters H135 Bearingless Main Rotor (BMR) blade Weber et al. (2021). The two datasets are selected due to their different levels of complexity in terms of system dimensions (1.45 m vs. 4.99 m span) and test conditions (shaker excitation vs. AVT). The identification accuracy of NExT-LF is evaluated against concurrent identifications obtained via SSI-CVA and NExT-ERA, as well as against available benchmark results. The overall objective is to assess whether the accuracy achieved by NExT-LF is suitable for aeronautical structures. To achieve this aim, the remainder of the paper is structured as follows:
@@
  • Section˜2 (Methodology) introduces the LF identification procedure, the output-only enabling method (NExT), and their coupling to reach NExT-LF;
@@ 2.1 Loewner Framework
The LF algorithm has previously been adopted for modelling multi-port electrical systems Lefteriu and Antoulas (2009), for aeroservoelastic modelling Quero et al. (2019), and for parametric model order reduction Vojkovic et al. (2023). Subsequently, the first and second authors employed LF to identify modal parameters in SIMO mechanical systems Dessena et al. (2023). More recent work has extended the LF to allow the extraction of modal parameters from multiple input, multiple output systems Dessena and Civera (2025); Dessena et al. (2026), output-only systems Dessena et al. (2025), and flutter estimation Dessena et al. (2025). The present study considers this latter output-only formulation.
To properly introduce LF, we first define the Loewner matrix $\bm{\mathbb{L}}$}: Let ($\mu_{j}$${v}_{j}$), j=1j=1,…,$q$ denote a row-wise collection of complex pairs and ($\lambda_{i}$${w}_{j}$), i=1i=1,…,$k$ a column-wise collection of complex pairs, with all points $\lambda_{i}$ and $\mu_{j}$ mutually distinct. The matrix $\bm{\bm{\mathbb{L}}}$, commonly termed the divided-differences matrix $\bm{\mathbb{L}}$}, is then given by:
@@
Karl Löwner first linked $\bm{\mathbb{L}}$ to rational interpolation, a setting that is frequently described as Cauchy interpolation Löwner (1934). Through this link, interpolants may be characterised via the determinants of suitable submatrices of $\bm{\mathbb{L}}$. Moreover, as shown in Antoulas et al. (2017); Mayo and Antoulas (2007), rational interpolants can be obtained directly from $\bm{\mathbb{L}}$. In the present work, the formulation adopted is the Loewner pencil, defined by the pair of matrices $\bm{\mathbb{L}}$ and $\bm{\mathbb{L}}_{s}$. The latter, $\bm{\mathbb{L}}_{s}$, indicates the shifted Loewner matrix.
@@
The Laplace-domain transfer function $\bm{H}(s)$ corresponding to $\bm{\Sigma}$ may be expressed as a rational matrix function $p\times m$, provided that the pencil $\bm{A}-\lambda\bm{E}$ is regular, i.e., $\bm{A}-\lambda\bm{E}$ is non-singular for at least one finite $\lambda\in\mathbb{C}$:
@@
Consider the general setting of tangential interpolation, which is commonly described as rational interpolation along tangential directions Kramer and Gugercin (2016). The associated right interpolation dataset is defined as:
@@
The values $\lambda_{i}$ and $\mu_{j}$ denote the sampling points at which $\bm{H}(s)$ is queried, which in the present context coincide with the chosen frequency bins. The vectors $\bm{r}_{i}$ and $\bm{l}_{j}$ specify the right and left tangential directions and are often drawn at random in practical implementations Quero et al. (2019). The quantities $\bm{w}_{i}$ and $\bm{v}_{j}$ are the corresponding tangential measurements. The rational interpolation task is then satisfied by enforcing the consistency between $\bm{w}_{i}$, $\bm{v}_{j}$ and the transfer function $\bm{H}$ associated with the realisation $\bm{\Sigma}$ in Equation˜2:
@@
A realisation of minimal order is attainable only when the underlying system is both controllable and observable. Under the assumption that the available samples originate from a system whose transfer function is given by Equation˜13, the generalised tangential observability matrix $\mathcal{O}_{v}$ and the generalised tangential controllability matrix $\mathcal{R}_{\rho}$ are introduced in Lefteriu and Antoulas (2010). Then it follows that Equation˜9 and Equation˜11 may be recast in the form:
@@
The expression in Equation˜20, which specifies the Loewner realisation in the presence of redundant data, is adopted as the reference formulation in the remainder of this work. The complete step-by-step procedure is provided in Mayo and Antoulas (2007); Antoulas et al. (2017). Finally, the modal parameters of the system can be obtained by performing an eigenanalysis of the matrices $\bm{A}$ and $\bm{C}$ that appear in Equation˜20. In practical modal analysis, this procedure is repeated over a range of model orders $k$, selected to capture the minimum set of modes of interest while identifying the highest order for which these modes remain stably estimated. At each instance, the frequency domain data, usually an FRF, $\bm{H}_{real}(s)$ is fed to the LF, which then breaks the data down, using the frequency bins as reference, into two disjoint subsets, the left and right data, so that the system matrices of the transfer function in Equation˜13 can be retrieved. It is worth noting that frequency band chosen must include the frequency of interest, as the interpolation of out-of-bounds data is not possible, i.e. all poles lay within the band examined.
@@ 2.2 Natural Excitation Technique
The NExT algorithm enables modal identification from output-only measurements by exploiting the statistical structure of ambient (unmeasured) excitations. In many operational settings, the applied forces cannot be reliably recorded (e.g., wind, traffic, or in-service aerodynamic loads), which prevents the direct construction of FRFs from measured inputs and outputs. NExT bypasses this limitation by assuming unknown excitation as a random process with a sufficiently broad frequency content to excite the modes of interest, and by using output correlations as impulse-response-type quantities III et al. (1993), the IRFs of the system can be obtained.
@@
where $\tau$ is the time delay. For finite records, (22) is estimated from sampled data. Using $N$ samples $y_{i}[k]$ and $y_{j}[k]$, a common unbiased estimator for non-negative lags is
@@
When the ambient excitation is broadband, and the system is LTI, the correlation functions can be written as a superposition of decaying sinusoids whose damped frequencies and decay rates coincide with those of the structural modes III et al. (1993). In practice, this means that $\widehat{R}_{y_{i}y_{j}}[\ell]$ (with $\ell$ denoting the discrete lag index, i.e. $\tau=\ell\Delta t$ for sampling period $\Delta t$) has the same functional form as a free-decay (IRF) for modal identification purposes, up to an unknown scaling and sensor-pair-dependent phase. A convenient representation is:
@@
where $n_{m}$ is the number of participating modes, $\omega_{r}$ and $\zeta_{r}$ are the natural frequency and damping ratio of the $r$-th mode, and $A_{r,ij}$ and $\theta_{r,ij}$ are constants that depend on the pair of sensors and the (unknown) spatial distribution of the excitation.
Given $p$ output channels, NExT forms a set of impulse-response-like sequences by selecting a reference channel $j$ and computing $\widehat{R}_{y_{i}y_{j}}[\ell]$ for $i=1,\dots,p$. The resulting collection $\{\widehat{R}_{y_{1}y_{j}},\widehat{R}_{y_{2}y_{j}},\dots,\widehat{R}_{y_{p}y_{j}}\}$ is interpreted as a multiple-output IRF associated with the chosen reference. An equivalent frequency-domain route, not considered in this work, computes cross-spectral densities and then obtains correlations via an inverse Fourier transform. Both approaches are consistent, provided that the stationarity assumptions are satisfied and the same normalisation conventions are used.